# A Generic Machine Learning Framework for Radio Frequency Fingerprinting

### v1

@@ 4.2 RFF-based Emitter Data Association
where ${\bf{g}}^{(a)}_{\phi}$ for this task is a distance metric (e.g. Euclidean, cosine distance, etc), ${\bf{F}}^{(a)}_{\theta}$ is a RFF function containing two separate projections of emitter data using the underlying RF fingerprint head ${\bf{f}}^{(a)}_{\theta}$ and ${\bf{x}}=({\bf{x}}_{1},{\bf{x}}_{2})\in{\mathcal{X}}^{(a)}\equiv\mathbb{R}^{d^{(a)}}\times\mathbb{R}^{d^{(a)}}$ where $d^{(a)}$ is the dimensionality of an emitter’s feature space for the EDA task. This results in the following optimization problem:
$\xi^{*}=\mbox{arg}\min_{\xi\in\Xi}\mathbb{E}_{({\bf{x}},y)\sim\mathcal{P}_{\mathcal{X}^{(a)}Y^{(a)}}}\mathcal{L}^{(a)}\left({\bf{h}}_{\xi}^{(a)}\left({\bf{x}}\right),y\right)$
where $\mathcal{P}_{\mathcal{X}^{(a)}{Y}^{(a)}}$ is the joint probability distribution of the RF feature space $\mathcal{X}^{(a)}$ and the task label space $\mathcal{Y}^{(a)}$ and $\mathcal{L}^{(a)}$ is the loss function for the EDA task.
Like the CS-SEI task, this optimization problem is solved in practice by collecting empirical datasets that are assumed to be sampled from the typically unknown joint probability distribution. Without loss of generality, it is assumed here that an empirical dataset $\mathcal{D}^{(a)}$ is transformed from the CS-SEI dataset $D^{(i)}$. In other words, $\mathcal{D}^{(a)}=\mathcal{M}(\mathcal{D}^{(i)})=\{({\bf{x}}_{1,k}^{(a)},{\bf{x}}_{2,k}^{(a)},y^{(a)}_{k})\}_{k=1}^{n_{a}}$, where $n_{a}$ is the number of samples in the dataset, $\mathcal{M}$ is a data transform operator such as the one discussed in Appendix A, $y^{(a)}_{k}\in\{0,1\}$ and $({\bf{x}}_{1,k}^{(a)},{\bf{x}}_{2,k}^{(a)})\in{\bf{X}}^{(a)}\equiv\mathbb{R}^{d^{(a)}}\times\mathbb{R}^{d^{(a)}}$. This results in the following optimization problem:
@@
In practice, for the data transformation method listed above, $n_{a}$ is selected according to a sampling strategy because for large $n_{i}$ the combinatorics become difficult to manage. We also note here that collecting EDA-specific empirical datasets is also possible.
4.3 Unsupervised Approach for RFEC
In this section, we formulate an unsupervised approach within the ML-based RFF framework. Typical methods for learning data structures and hidden patterns without labels are clustering approaches such as DBSCAN [12] or dimensionality reduction techniques (e.g. Principal Component Analysis (PCA) followed by some method of clustering). Here we describe one which uses an auto-encoder approach for RFEC. Autoencoders typically involve projecting the original data into a lower dimensional representation space, where the encoded data is then fed through a decoder to reconstruct the original input. If the reconstruction error is small, then the encoded representation is assumed to contain important information of the original data. It is a dimensionality reduction technique similar to PCA. In this setting, the optimization problem is:
@@ 4.3 Unsupervised Approach for RFEC
where ${\bf{\hat{x}}}={\bf{g}}^{(r)}_{\phi}({\bf{F}}^{(r)}_{\theta}({\bf{x}}))$, $\mathcal{P}_{{\mathcal{X}}^{(r)}}$ is the probability distribution of the emitter feature space $\mathcal{X}^{(r)}$, $\mathcal{L}^{(r)}$ is the reconstruction loss associated with the auto-encoding approach, ${\bf{g}}^{(r)}_{\phi}$ in this task is the decoder and ${\bf{F}}^{(r)}_{\theta}\equiv{\bf{f}}_{\theta}$ is the projection of the data into the RFF representation space.
In practice, this optimization problem is solved by collecting an empirical dataset that is assumed to be sampled from $\mathcal{P}_{\mathcal{X}^{(r)}}$ IID $n_{r}$ times, obtaining a dataset $\mathcal{D}^{(r)}=\{({\bf{x}}^{(r)}_{k})\}_{k=1}^{n_{r}}$, where ${\bf{x}}^{(r)}_{k}\in{\mathcal{\bf{X}}}^{(r)}$ with ${\mathcal{\bf{X}}}^{(r)}$ being the sample dataset. Then, the optimization problem is:
@@
where $\hat{{\bf{x}}}_{k}={\bf{g}}^{(r)}_{\phi}\left({\bf{F}}^{(r)}_{\theta}\left({\bf{x}}_{k}\right)\right)$.
4.4 Open Set SEI and EDA
@@
A key challenge when solving any of the above tasks is generalization towards emitter data not contained in the training dataset. Open Set Recognition (OSR) methods determine whether a sample belongs to the original training distribution or not. In the CS-SEI regime, a finite number of emitters are assumed in existence and a model is trained to discriminate between them; predicated on the Closed World Assumption (CWA) [14]. From the OSR lens, the key challenge following training in the CS-SEI regime is to determine whether an individual sample presented to the trained model is an observation from a known emitter or an unknown one. The CWA ensures false positives occur on open set data since the model is designed to predict an emitter class from the set of known emitters. In contrast, the key Open Set (OS) challenge in EDA is understanding when the comparator has observed enough emitters in training such that it generalizes well to unknown ones. In other words, EDA learns to compare and is not conditioned on the emitter class. Without loss of generality, we introduce the OSR-SEI problem within the RFF ML framework using some terminology from [56].
@@
Let us assume two sample datasets $({\bf{X}}_{known}^{(i)},{\bf{Y}}_{known}^{(i)})$ and $({\bf{X}}_{unknown}^{(i)},{\bf{Y}}_{unknown}^{(i)})$ are drawn from the joint probability distribution $\mathcal{P}_{\mathcal{X}^{(i)}\mathcal{Y}^{(i)}}$ such that ${\bf{Y}}^{(i)}_{known}\cap{\bf{Y}}^{(i)}_{unknown}$$\emptyset$, denoting ${\bf{Y}}^{(i)}_{known}$ as the label set of known emitters to the CS-SEI model and ${\bf{Y}}^{(i)}_{unknown}$ as the label set of unknown emitters to the CS-SEI model. Then, we define the CS-SEI dataset $\mathcal{D}_{c}^{(i)}=\left\{({\bf{x}}_{k}^{(i)},y_{k}^{(i)})\right\}_{k=1}^{n_{i}}$, where ${\bf{x}}_{k}^{(i)}\in{\bf{X}}^{(i)}_{known}$, $y_{k}\in{\bf{Y}}^{(i)}_{known}$, and $n_{i}$ is the total number of samples. In addition, we define a OSR-SEI dataset $\mathcal{D}_{o}^{(i)}=\{({\bf{x}}_{k}^{(i)},y^{(i)}_{k})\}_{k=1}^{n_{o}}\subset({\bf{X}}_{known}^{(i)}\bigcup{\bf{X}}_{unknown}^{(i)})\times({\bf{Y}}^{(i)}_{known}\bigcup{\bf{Y}}^{(i)}_{unknown})$ which may or may not contain unknown unique emitters, with $n_{o}$ being the total number of samples. Then, we combine the two into a dataset $\mathcal{D}^{(osr,i)}=\{\mathcal{D}_{c}^{(i)},\mathcal{D}_{o}^{(i)}\}$.
@@
In the OSR setting, the goal is to simultaneously learn a model capable of solving the CS-SEI problem defined in Section 4.1 as well as produce a score indicating whether a sample belongs to $\it{any}$ of the known classes. We define a score function $S(x)\in\{0,1\}$, which indicates whether a sample $x$ belongs to the class of known emitters or the class of unknown emitters (0 indicates $x$ belonging to a known class whereas $1$ indicates $x$ belonging to an unknown class). The two most common ways $S$ can be computed are:
@@
    Jointly learn $S(x)$ by using $\mathcal{D}_{c}^{(i)}$ for CS-SEI model training and $\mathcal{D}^{(osr,i)}=\{\mathcal{D}_{c}^{(i)},\tilde{\mathcal{D}}_{o}^{(i)}\}$ for binary classification between known and unknown emitters (where $\tilde{\mathcal{D}}_{o}^{(i)}$ is a proportion of $\mathcal{D}_{o}^{(i)}$) then evaluate performance on remaining OSR data $\mathcal{D}_{o}^{(i)}\backslash\tilde{\mathcal{D}}_{o}^{(i)}$;
@@
    Conduct CS-SEI model training using $\mathcal{D}_{c}^{(i)}$ then use rule-based strategies to compute $S(x)$ (e.g. Maximum Softmax Probability (MSP) with threshold [56] [48] ) and use $\mathcal{D}_{o}^{(i)}$ to assess OSR performance.
A hybrid of the above methods can also be applied (e.g. ensemble methods).
Within the generic RFF framework, either of these tasks translate as solving the following optimization problem:
@@
where ${\bf{h}}^{(osr,i)}_{\xi}({\bf{x}})={\bf{g}}^{(osr,i)}_{\phi}\left({\bf{F}}^{(osr,i)}_{\theta}\left({\bf{x}}\right)\right)$, ${\bf{x}}=({\bf{x}}_{c},{\bf{x}}_{o})$ and ${\bf{y}}=(y_{c},y_{o})$ with $({\bf{x}}_{c},y_{c})\in\mathcal{D}_{c}^{(i)}$ and $({\bf{x}}_{o},y_{o})\in\mathcal{D}_{o}^{(i)}$. In this case, ${\bf{g}}_{\phi}^{(osr,i)}$ is not only the classifier head but also the open set score computer and $\mathcal{L}^{(osr,i)}$ could be an OSR-specific objective function such as that in [6]. Once a methodology for scoring has been determined and solved, it can then be used to conduct incremental learning [41] [27], perform CS-SEI with quasi-labelling of OSR samples [28], amongst other methods, to address the OS problem. Figure 2 shows an example pipeline for incorporating OS-SEI.
@@
5 Experimentation
This section provides a description of real datasets and model architectures used for experimentation as well as documenting experiments which solve a selection of tasks from Section 4.
5.1 Datasets
5.1.1 Digital Mobile Radios: SIGINT
@@
The Digital Mobile Radio (DMR) dataset was collected from eight Commercial Off-The-Shelf (COTS) handsets using a Lime Software Defined Radio (LimeSDR) in four environments; an open field with direct line-of-sight between each DMR and the SDR, two semi-urban environments and one urban environment. The DMR devices are from five different manufacturers, namely one Hytera PD705, one TYT, two Kenwood NX1300D (i.e. identical models), two Entel DX482, and two Motorola DP4400E. All DMRs were programmed to operate in the 433 MHz band where they can use one of the 14 channels, each having a bandwidth of approximately 12.5 kHz. The sampling rate at the receiver was set at approximately 200 kHz. As part of the data pre-processing steps, a burst segmentation algorithm was applied for segments of 30 seconds of continuous transmission to isolate individual bursts from each DMR handset. RFF and identification of individual DMR handsets has several applications in SIGINT and ISR contexts such as determining presence of emitters of interest (e.g. associated with high value targets or threats), tracking them over time and/or building pattern-of-life of their operators.
5.1.2 AIS: Spaceborne Maritime Surveillance
@@
The Automatic Identification System (AIS) dataset contains transmissions from vessels in the Panama Canal area as captured by SDRs on Low Earth Orbit (LEO) satellites. This proprietary dataset comprises of approximately 10 minutes of RF recording of the VHF channels dedicated to ground maritime AIS emitters. Bespoke pre-processing algorithms are applied to detect-extract the individual AIS bursts and, where possible, demodulate-decode them to obtain their unique Maritime Mobile Service Identity (MMSI) number, which can serve as a label. It is noted that classical demodulating and decoding techniques for AIS often require relatively high-quality AIS transmissions, for instance a substantially high SNR and/or low levels of interference. The AIS dataset utilised here has over 600 AIS bursts from 48 sources such that each emitter (i.e. with a unique MMSI) has at least 10 transmissions. The sampling rate is approximately 38.4 kHz and the AIS burst time slot is 26.67 m/s.
RFF-based EDA can improve geolocating the transmission source, since by identifying various transmissions that originate from the same source, we can effectively constrain the observables (time and frequency difference of arrival measurements) and dwell on the emitter over a much longer time intervals, thereby utilising multiple emissions in the estimation of the emitters’ locations. Whilst geolocation of ground RF emitters from LEO satellite clusters and subsequently tracking changes in their positions (including from multiple passes) has many ISR applications, spaceborne RFF facilitates delivering enhanced situational awareness. For example, results from SEI or EDA can be fused with other data sources, such as images, to confirm identify of vessels and/or detect malicious activities (e.g. spoofing).
5.1.3 Data from Drones: Counter-UAS

### v2

@@ IV-B RFF-based Emitter Data Association
where ${\bf{g}}^{(a)}_{\phi}$ here is a distance metric (e.g. Euclidean, cosine distance, etc.), ${\bf{F}}^{(a)}_{\theta}$ is a RFF function containing two separate projections of emitter data using the underlying RF fingerprint head ${\bf{f}}^{(a)}_{\theta}$ and ${\bf{x}}=({\bf{x}}_{1},{\bf{x}}_{2})\in{\mathcal{X}}^{(a)}\equiv\mathbb{R}^{d^{(a)}}\times\mathbb{R}^{d^{(a)}}$ where $d^{(a)}$ is the dimensionality of an emitter’s feature space for the EDA task. This results in the following optimisation problem:
$\xi^{*}=\mbox{arg}\min_{\xi\in\Xi}\mathbb{E}_{({\bf{x}},y)\sim\mathcal{P}_{\mathcal{X}^{(a)}Y^{(a)}}}\mathcal{L}^{(a)}\left({\bf{h}}_{\xi}^{(a)}\left({\bf{x}}\right),y\right),$
where $\mathcal{P}_{\mathcal{X}^{(a)}{Y}^{(a)}}$ is the joint distribution of the RF feature space $\mathcal{X}^{(a)}$ and the task label space $\mathcal{Y}^{(a)}$; $\mathcal{L}^{(a)}$ is the EDA loss function.
Like CS-SEI, this optimisation problem is solved in practice by collecting empirical datasets, assumed to be sampled from the typically unknown joint probability distribution. Without loss of generality, it is assumed here that an empirical dataset $\mathcal{D}^{(a)}$ is transformed from the CS-SEI dataset $D^{(i)}$. In other words, $\mathcal{D}^{(a)}=\mathcal{M}(\mathcal{D}^{(i)})=\{({\bf{x}}_{1,k}^{(a)},{\bf{x}}_{2,k}^{(a)},y^{(a)}_{k})\}_{k=1}^{n_{a}}$, where $n_{a}$ is the number of samples in the dataset, $\mathcal{M}$ is a data transform operator such as the one outlined in Appendix A, $y^{(a)}_{k}\in\{0,1\}$ and $({\bf{x}}_{1,k}^{(a)},{\bf{x}}_{2,k}^{(a)})\in{\bf{X}}^{(a)}\equiv\mathbb{R}^{d^{(a)}}\times\mathbb{R}^{d^{(a)}}$. This results in:
@@
For the data transformation method presented in Appendix A, $n_{a}$ is selected according to a sampling strategy because for large $n_{i}$ the combinatorics become difficult to manage. We also note here that collecting EDA-specific empirical datasets is also possible.
IV-C Unsupervised Learning for RFEC
We recall that RFEC aims to determine and group the received RF transmissions based on the their sources, where the number and identity of the encountered emitters are unknown a priori. Hence no labels are available and unsupervised learning methods are applied. The obtained “clustering“ and distinctions between the observed groups of emitters is expected to be driven by the automatic extraction of RF fingerprints [62, 1, 2]. Common methods for learning data structures and hidden patterns without labels are clustering approaches such as, K-means, Gaussian mixture models, hierarchical clustering, DBSCAN [63] and many others. These can be applied on the output of a dimensionality reduction technique, such as principal component analysis (PCA), or low-dimensional representations of the data inclusive within an auto-encoder architecture [33]. The latter is adopted here where we describe an AE-based RFEC.
Autoencoders typically involve projecting the original data into a lower dimensional representation space, where the encoded data is then fed through a decoder to reconstruct the original input. If the reconstruction error is small, then the encoded representation is assumed to contain the important salient information of the original data. In this setting, the optimisation problem is:
@@ IV-C Unsupervised Learning for RFEC
where ${\bf{\hat{x}}}={\bf{g}}^{(r)}_{\phi}({\bf{F}}^{(r)}_{\theta}({\bf{x}}))$, $\mathcal{P}_{{\mathcal{X}}^{(r)}}$ is the probability distribution of the emitter feature space $\mathcal{X}^{(r)}$, $\mathcal{L}^{(r)}$ is the reconstruction loss associated with the auto-encoding approach, ${\bf{g}}^{(r)}_{\phi}$ in this task is the decoder and ${\bf{F}}^{(r)}_{\theta}\equiv{\bf{f}}_{\theta}$ in the context of the RFF framework is viewed as the projection of the data into the emitter’s RF fingerprint representation space.
In practice, this optimisation problem is solved by collecting an empirical dataset that is assumed to be sampled from $\mathcal{P}_{\mathcal{X}^{(r)}}$ IID $n_{r}$ times, resulting in a dataset $\mathcal{D}^{(r)}=\{({\bf{x}}^{(r)}_{k})\}_{k=1}^{n_{r}}$, where ${\bf{x}}^{(r)}_{k}\in{\mathcal{\bf{X}}}^{(r)}$ with ${\mathcal{\bf{X}}}^{(r)}$ being the sample dataset. This leads to the following optimisation problem is:
@@
where $\hat{{\bf{x}}}_{k}={\bf{g}}^{(r)}_{{\phi}^{*}}\left({\bf{F}}^{(r)}_{\theta^{*}}\left({\bf{x}}_{k}\right)\right)$. To group the processed RF emissions, any clustering algorithm can be applied to the learnt signal representation (i.e. fingerprint as in Figure 1) in a lower dimensional space, this is ${\hat{\bf{z}}}_{k}={\bf{F}}^{(r)}_{\theta^{*}}\left({\bf{x}}_{m}\right)$ for a given input $\bf{x}_{m}$, based on the encoder output; see Section VI-E for an illustration of RFEC on real RF captures.
IV-D Open Set SEI and EDA
@@
When solving any of the above RFF-dependent tasks, the generalisation of the ML model to emitter data not contained in the training dataset could be a fundamental challenge. Open Set Recognition (OSR) methods typically first determine whether a sample belongs (or is sufficiently close) to the original training data distribution or not. In the CS-SEI regime, a finite number of emitters are assumed in existence and a model is trained to discriminate between them; predicated on the closed world assumption (CWA) [64]. From the OSR lens, the key difficulty following training in the CS-SEI regime is to establish whether an individual sample presented to the trained model is an observation from a known emitter or an unknown one. The CWA ensures false positives occur on open set data since the model is designed to predict an emitter class from the set of known emitters. In contrast, the critical open set (OS) obstacle in EDA is understanding when the comparator has observed enough emitters in training such that it generalises well to unknown ones. In other words, EDA learns to compare and is no.t conditioned on the emitter class. Without loss of generality, we introduce the OSR-SEI problem within the framework in Sections III and IV using terminology from [65].
@@
Let us assume two sample datasets $({\bf{X}}_{known}^{(i)},{\bf{Y}}_{known}^{(i)})$ and $({\bf{X}}_{unknown}^{(i)},{\bf{Y}}_{unknown}^{(i)})$ are drawn from the joint probability distribution $\mathcal{P}_{\mathcal{X}^{(i)}\mathcal{Y}^{(i)}}$ such that ${\bf{Y}}^{(i)}_{known}\cap{\bf{Y}}^{(i)}_{unknown}$$\emptyset$. ${\bf{Y}}^{(i)}_{known}$ and ${\bf{Y}}^{(i)}_{unknown}$ denote the label set of known and unknown emitters to the CS-SEI model, respectively. Then, we define the CS-SEI dataset $\mathcal{D}_{c}^{(i)}=\left\{({\bf{x}}_{k}^{(i)},y_{k}^{(i)})\right\}_{k=1}^{n_{i}}$, where ${\bf{x}}_{k}^{(i)}\in{\bf{X}}^{(i)}_{known}$, $y_{k}\in{\bf{Y}}^{(i)}_{known}$, and $n_{i}$ is the total number of samples. In addition, we define a OSR-SEI dataset $\mathcal{D}_{o}^{(i)}=\{({\bf{x}}_{k}^{(i)},y^{(i)}_{k})\}_{k=1}^{n_{o}}\subset({\bf{X}}_{known}^{(i)}\bigcup{\bf{X}}_{unknown}^{(i)})\times({\bf{Y}}^{(i)}_{known}\bigcup{\bf{Y}}^{(i)}_{unknown})$ which may or may not contain unknown unique emitters, with $n_{o}$ being the total number of samples. We can combine both into a single dataset as per $\mathcal{D}^{(osr,i)}=\{\mathcal{D}_{c}^{(i)},\mathcal{D}_{o}^{(i)}\}$.
@@
In the OSR setting, the goal is to simultaneously learn a model capable of solving the CS-SEI problem defined in Section IV-A and produce a score indicating whether a sample belongs to $\it{any}$ of the known classes. We define a score function $S(x)\in\{0,1\}$, which indicates whether a sample $x$ belongs to the class of known emitters or the class of unknown emitters (0 indicates $x$ belonging to a known class and 1 otherwise). The two most common approaches to determining this score are:
@@
    Jointly learn $S(x)$ by using $\mathcal{D}_{c}^{(i)}$ for CS-SEI model training and $\mathcal{D}^{(osr,i)}=\{\mathcal{D}_{c}^{(i)},\tilde{\mathcal{D}}_{o}^{(i)}\}$ for binary classification between known and unknown emitters, where $\tilde{\mathcal{D}}_{o}^{(i)}$ is a proportion of $\mathcal{D}_{o}^{(i)}$, then evaluate performance on remaining OSR data $\mathcal{D}_{o}^{(i)}\backslash\tilde{\mathcal{D}}_{o}^{(i)}$, and;
@@
    Conduct CS-SEI model training using $\mathcal{D}_{c}^{(i)}$ followed by applying rule-based strategies to compute $S(x)$, for example via maximum softmax probability (MSP) with threshold [65] [66] and then employing $\mathcal{D}_{o}^{(i)}$ to assess the OSR performance.
Whilst a hybrid of the two techniques can be utilised (e.g. ensemble methods), within the presented generic RFF framework either correspond to solving the following optimisation problem:
@@
where ${\bf{h}}^{(osr,i)}_{\xi}({\bf{x}})={\bf{g}}^{(osr,i)}_{\phi}\left({\bf{F}}^{(osr,i)}_{\theta}\left({\bf{x}}\right)\right)$, ${\bf{x}}=({\bf{x}}_{c},{\bf{x}}_{o})$ and ${\bf{y}}=(y_{c},y_{o})$ with $({\bf{x}}_{c},y_{c})\in\mathcal{D}_{c}^{(i)}$ and $({\bf{x}}_{o},y_{o})\in\mathcal{D}_{o}^{(i)}$. In this case, ${\bf{g}}_{\phi}^{(osr,i)}$ is not only the classifier head but also the open set score computer and $\mathcal{L}^{(osr,i)}$ could be an OSR-specific objective function such as that in [67]. Once a methodology for scoring has been established and solved, it can then be used to conduct incremental learning [68] [27], perform CS-SEI with quasi-labeling of OSR samples [43], using outlier exposure with label smoothing [23], amongst other methods, to address the OS problem. An alternative is meta-learning type techniques which can be leveraged for few-shot open-set recognition as in [42]. Figure 2 depicts an example pipeline for incorporating OS-SEI. It is underpinned by: a) a suitable out-of-distribution detector (OOD) [69] to reveal out of distribution samples (i.e. those that belong to a emitter not currently in the closed-set list of classes); and b) retraining, including incrementally, the CS-SEI model to address the detected sample(s) from new emitters and their corresponding (quasi-)labels, where relevant.
@@
V Experimentation: Datasets, Models and Set-up
In this section, we first describe the real RF datasets, the ML model architectures and the general experimentation set for a number of disparate RFF-dependent tasks.
V-A Datasets
V-A1 Digital Mobile Radios: SIGINT
@@
The DMR dataset was collected from eight commercial off-the-shelf (COTS) handsets using the COTS software defined radio (SDR) LimeSDR in four environments; an open field with direct line-of-sight between each DMR and the SDR, two semi-urban environments and one urban environment. The DMR devices are from five different manufacturers, namely one Hytera PD705, one TYT, two Kenwood NX1300D (i.e. identical models), two Entel DX482, and two Motorola DP4400E. All DMRs were programmed to operate in the 433 MHz band where they can use one of 14 channels, each with a bandwidth of approximately 12.5 kHz. The sampling rate at the receiver was set at approximately 200 kHz. As part of the data pre-processing steps, a burst segmentation algorithm was applied for segments of 30 seconds of continuous transmission to isolate individual bursts from each DMR handset. RFF and identification of individual DMRs has several applications in SIGINT and ISR contexts such as determining the presence of emitters of interest (e.g. associated with high value targets or threats), tracking them over time and/or building pattern-of-life of their operators.
V-A2 AIS: Spaceborne Maritime Surveillance
@@
The automatic identification system dataset contains transmissions from vessels in and near the Panama Canal area as captured by SDRs on Low Earth Orbit (LEO) satellites. This proprietary dataset comprises of approximately 10 minutes of RF recordings of the VHF channels dedicated to ground maritime AIS emitters. Bespoke pre-processing algorithms are applied to detect-extract the individual AIS bursts and, where possible, demodulate-decode them to obtain their unique maritime mobile service identity (MMSI) number, which can serve as a label. It is noted that classical demodulating-decoding techniques for AIS often require relatively high-quality AIS transmissions, for instance a substantially high SNR and/or low levels of interference. The AIS dataset utilised here has over 600 AIS bursts from 48 sources such that each emitter has at least 10 detected transmissions. The sampling rate is approximately 38.4 kHz with an AIS burst time slot of 26.67ms. RFF-based EDA can improve geolocating the transmission source. By identifying various transmissions that originate from the same source, we can effectively constrain the observables (time and frequency difference of arrival measurements) and dwell on the emitter over a much longer time intervals. Thereby, we can utilise multiple emissions in the estimation of the emitters’ locations. Whilst geolocation of ground RF emitters from LEO satellite clusters and subsequently tracking changes in their positions (including from multiple passes) has many ISR applications, spaceborne RFF facilitates delivering enhanced situational awareness. For example, results from SEI or EDA can be fused with other data sources, such as images, to confirm the identity of vessels and/or detect malicious activities (e.g. AIS spoofing).
V-A3 RF Data from Drones: Counter-UAS